A fine foods store sells excellent Mozart Cheese and sells an average of 64 pounds per month with a known standard deviation of 10 pounds. If the manager would like to meet demand 90% of the time, how many pounds should she start the month with? 76.8 74.2 72.6 71.1
Answer #1

μ = 64, σ = 10

z- score such that the area under the standard normal curve to its right is 90% is z = 1.2816

x = μ + z * σ = 64 + 1.2816 * 10 = 76.82 pounds

Answer: Option (a)